A glass sphere is filled to full volume with a gas. The pressure of the gas inside the sphere is 30.0 atm, and the temperature is 30.0°C. The sphere is taken outside on a cold day. The temperature of the gas decreases to 10.0°C. What is the new pressure of the gas? Assume that the volume is constant.


1
Expert's answer
2021-02-11T04:29:00-0500

P1 = 30.0 atm

T1 = 30.0°C + 273 = 303 K

V1 = V2 = const

P2 = unknown

T2 = 10.0°C + 273 = 283 K


Solution:

Since the volume doesn't change, Gay-Lussac's Law can be used.

Gay-Lussac's Law can be expressed as: P1/T1 = P2/T2

P1T2 = P2T1

To find the final pressure, solve the equation for P2:

P2 = P1T2/T1

P2 = (30.0 atm × 283 K) / (303 K) = 28.0198 atm = 28.0 atm

P2 = 28.0 atm


Answer: 28.0 atm is the new pressure of the gas.